OpenStudy (anonymous):

Solve the equation for x: 2x - 16y = 10

Parth (parthkohli):

A good start is adding 16y to both sides. \( \color{Black}{\Rightarrow 2x = 10 + 16y}\) No need to divide both sides by 2 in the old fashioned way, but you can factor. \( \color{Black}{\Rightarrow 2(x) = 2(5 + 8y) }\) Now, cancel those 2's. \( \color{Black}{\Rightarrow x = 5 + 8y}\)